OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). A. y + 3 = -5(x - 2) B. y - 2 = -5(x + 3) C. y + 3 = -1/5 (x - 2) D. y - 2 = -1/5 (x + 3)

OpenStudy (anonymous):

\[\Large y-y_1=\frac{y_2-y_1}{x_2-x_1}\cdot (x-x_1)\] just substitute...

OpenStudy (anonymous):

agree